A 20 g block of ice is cooled to −65 ◦C. It is added to 577 g of water in an 80 g copper calorimeter at a temperature of 29◦C. Find the final temperature. The specific heat of copper is 387 J/kg · ◦C and of ice is 2090 J/kg · ◦C . The latent heat of fusion of water is 3.33 × 10^5 J/kg and its specific heat is 4186 J/kg · ◦C . Answer in units of ◦C.


1
Expert's answer
2021-05-27T17:48:29-0400

Specific heat of copper is "C_{Cu}=387\ J/(kg\cdot\degree C)"

specific heat of water is "C_{w}=4186\ J/(kg\cdot\degree C)"

specific heat of ice is "C_{ice}=2100\ J/(kg\cdot\degree C)"

he latent heat of fusion of water is "\lambda=3.33⋅10 ^5\ J/kg"


The heat capacity of water seams to be much greater then of the ice, thus we will decide the ice will melt and final temperature of system "T>0\degree C". Determine the heat capacity wich must be added to block of ice. This consist of three terms. First is the heat capacity to warm ice block to "0\degree C".

"Q_{ice1}=C_{ice}m_{ice}\Delta T"

"m_{ice}=0.020\ kg"

"\Delta T=0-(-65)=65\ \degree C"

"Q_{ice1}=2100\cdot0.02\cdot65=2730\ J"


Second is the heat to melt ice block at "0\ \degree C" :

"Q_{ice2}=\lambda m_{ice}=3.33\cdot10^5\cdot0.02=6660\ J"


Third is to warm water of ice block to final temperature:

"Q_{ice3}=C_wm_{ice}T"

The law of conservation of thermal energy in this case says that all this energy (1)-(3) ice gets from cooled water and copper calorimeter. The water lost the amount of heat 

"Q_{w}=C_wm_{w}(T-T_{init})"

"m_{w}=0.577\ kg" , "T_{init}=29\ \degree C"

The heat of copper changed by

"Q_{w}=C_{Cu}m_{Cu}(T-T_{init})"

"m_{Cu}=0.080\ kg"


The law of conservation of energy in calorimeter is

"Q_{ice1}\u200b +Q_{ice2}\u200b +Q_{ ice3}\u200b +Q_{ w}\u200b +Q_{ Cu}\u200b =0"


Then the final temperature of system is


"T=\frac{(C_wm_w+C_{Cu}m_{Cu})T_{init}-Q_{ice1}-Q_{ice2}}{C_wm_w+C_{Cu}m_{Cu}+C_{ice}m_{ice}}=\frac{29(4186\cdot0.577+387\cdot0.08)-2730-6660}{4186\cdot0.577+387\cdot0.08+2100\cdot0.02}=24.7\ \degree C"